# On Variants of Inverse Cluster Size Problem & General Magnification

### v1

@@ Abstract.
In this article we establish certain variants of the Inverse Cluster Size problem. We introduce the notion of primitive extensions and establish the Primitive variant of the problem. Precisely, we prove the existence of primitive extensions over number fields of any given degree and cluster size less than the degree. We also introduce the notions of Strong and Weak General Magnification and the notion of general primitive extensions. We establish some interesting cases of the General primitive variant of the problem.
@@ 1. Introduction
The Inverse Cluster Size Problem for Number Fields was established by the first author and Bhagwat in [Bhagwat_2025] generalizing a result of Perlis in [perlis2004roots] for number fields and improving on the generalization proved previously by the second and third author in [krithika2023root]. In this article we establish certain variants of the problem.
Let $K$ be a perfect field. We fix an algebraic closure $\bar{K}$ once and for all and work with finite extensions of $K$ contained in $\bar{K}$. Let $L/K$ be a degree $n$ extension and $\tilde{L}$ be its Galois closure inside $\bar{K}$. Let $G={{\rm Gal}}(\tilde{L}/K)$ and $H={{\rm Gal}}(\tilde{L}/L)$. We have the notion of cluster size of $L/K$, $r_{K}(L)$ which is $N_{G}(H):H]$ (See Section 2.1 in [Bhagwat_2025] for basic properties of cluster size of field extension). From Section 3.2 in [Bhagwat_2025], number of clusters of $L/K$, $s_{K}(L)$ is $G:N_{G}(H)]$ which is also the number of distinct fields inside $\bar{K}$ isomorphic to $L$ over $K$.
In Section 2, the notion of an extension being primitive is introduced. We recall the notions of unique descending chains and unique ascending chains for extensions introduced in [Bhagwat_2025]. We then establish a criterion in Theorem 2.4 for an extension to be primitive in terms of unique chains which serves as a major ingredient in the proof of our Theorem 2.6 which is as follows.
@@
In Section 3, the notions of Strong and Weak General Magnification are introduced and studied inspired from the notions of Strong and Weak Cluster Magnification in [Bhagwat_2025] and motivated from results in [krithika2025inflated]. In Section 4, the notion of an extension being general primitive is introduced. We then go on to establish some interesting cases of the General Primitive Inverse Cluster Size Problem which is encapsulated in the following result.
@@ 2.1. Primitive Extensions
The following is reformulation of Definition 4.2.4 in [Bhagwat_2025] for field extensions (This notion for polynomials occurs in [krithika2023root] as well).
@@
Consider $L/K$. Let $G=Gal(\tilde{L}/K)$ and $H=Gal(\tilde{L}/L)$. If $H$ is not contained in any proper normal subgroup of $G$ then $L/K$ is primitive. In particular, if $G$ is simple then $L/K$ is primitive.
@@ 2.2. Unique Chains for Extensions and Primitive Extensions
We have the notions of unique descending chains and unique ascending chains for extensions introduced in Section 7 in [Bhagwat_2025]. We recall those notions here. Let $L/K$ be a nontrivial finite extension.
@@
Let $\tilde{L}$ be Galois closure of $L/K$ and $G=Gal(\tilde{L}/K)$ and $H=Gal(\tilde{L}/L)$. By Theorem 7.1.1 in [Bhagwat_2025], $N_{1}=\tilde{L}^{N_{G}(H)}$. Now since $L=L^{\prime}F$ and $\tilde{L^{\prime}}\cap F=K$, we have that $L/L^{\prime}$ is Galois. Hence $L^{\prime}=\tilde{L}^{H^{\prime}}$ where $H\unlhd H^{\prime}\subset G$. Thus $H^{\prime}\subset N_{G}(H)$. Hence $N_{1}\subset L^{\prime}$. Thus $N_{i}\subset L^{\prime}$ and hence $F\subset L^{\prime}$. This forces $F=K$ which gives a contradiction to the magnification being nontrivial. ∎
@@
A shorter proof of above theorem by using results from Section 8.4 in [Bhagwat_2025]: We have $N_{i}=F_{j}\neq L,K$. Suppose $L/K$ is obtained by nontrivial strong cluster magnification from some subextension say $L^{\prime}/K$ through $F/K$. By Theorem 8.4.1 and Remark 8.4.2 in [Bhagwat_2025], $N_{i}\subset L^{\prime}$. By Theorem 8.4.3 and Remark 8.4.4 in [Bhagwat_2025], $F\subset F_{j}$. Thus $F\subset L^{\prime}$ which gives a contradiction.
@@
The assumption $N_{i}=F_{j}\neq L,K$ is important in the theorem. Suppose $L/K$ is Galois with Galois group $A\times B$ for $A$ and $B$ nontrivial groups and $|A|>2$. The unique descending chain for $L/K$ is $L=N_{0}\supsetneq K=N_{1}$ and unique ascending chain for $L/K$ is $K=F_{0}\subsetneq L=F_{1}$. Thus $N_{0}=F_{1}=L$ and $N_{1}=F_{0}=K$. But by Corollary 4.1.8 in [Bhagwat_2025] we have that $L/K$ is not primitive.
@@
The converse of Theorem 2.4 is false. Let $f$ over $K$ be irreducible of deg $n>2$ with Galois group ${\mathfrak{S}}_{n}$ with roots $\alpha_{i}\in\bar{K}$ for $1\leq i\leq n$. For $1\leq k\leq n-2$, let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$. Assume $k\neq 1,n/2$. By Theorem 7.3.6 in [Bhagwat_2025], the unique descending chain for $L_{k}/K$ is $L_{k}\supsetneq N_{k}$ and unique ascending chain for $L_{k}/K$ is $K$. Hence $L_{k}/K$ doesn’t satisfy hypothesis of Theorem 2.4 but we have that $L_{k}/K$ is primitive. This is because we don’t have any $K\subsetneq F\subset L_{k}$ such that $F/K$ is Galois.
@@
Since $c$ is odd, it does not divide the discriminant of $\mathbb{Q}(\zeta)/\mathbb{Q}$ where $\zeta$ is a primitive $n$-th root of unity in $\bar{\mathbb{Q}}$. Thus $\sqrt{c}\not\in\mathbb{Q}(\zeta)$. By Proposition 7.3.1 in [Bhagwat_2025], the unique descending chain is $L=N_{0}\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}=K$ and the unique ascending chain is $K=F_{0}\subsetneq F_{1}\subsetneq\dots\subsetneq F_{k}=L$ with $N_{i}=F_{k-i}=\mathbb{Q}(a^{2^{i}})$ for all $0\leq i\leq k$. Since $k\geq 2$, we are done by Theorem 2.4.
@@ 2.3. Primitive Inverse Cluster Size Problem
We proceed in the same way as the first author and Bhagwat proceeded in the proof of Inverse Cluster Size Problem for Number Fields Theorem 3.1.1 in [Bhagwat_2025].
For $r=1$ we have that there exists $L/K$ of degree $n$ with $Gal(\tilde{L}/K)$ as ${\mathfrak{S}}_{n}$. This $L/K$ satisfies $r_{K}(L)=1$. The subgroup of ${\mathfrak{S}}_{n}$ fixing $L$ is isomorphic to ${\mathfrak{S}}_{n-1}$. Since ${\mathfrak{S}}_{n-1}$ is not contained in any proper normal subgroup of ${\mathfrak{S}}_{n}$. Hence by Proposition 2.3, $L/K$ is primitive.
For $r>1$ we have that there exists $L/K$ with $G=Gal(\tilde{L}/K)$ solvable such that action of the group is transitive on $n$ points, and a point stabiliser fixes precisely $r$ points. We have $G=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}}$ with semidirect product group law
@@
Since $1<r<n$. Thus by Theorem 7.3.4 in [Bhagwat_2025], the unique descending chain is $L\supsetneq N\supsetneq K$ and the unique ascending chain is $K\subsetneq F\subsetneq L$ and they both coincide. Hence $N=F\neq L,K$. Thus by Theorem 2.4, $L/K$ is primitive. ∎
@@
The same proof works for the case $r=n=p^{k}>2$ where $p$ is a prime. In that case $L/K$ is Galois with $Gal(L/K)$ as $\mathbb{Z}/p^{k}\mathbb{Z}$. By uniqueness in structure theorem for finite abelian groups, $Gal(L/K)$ is not isomorphic to $A\times B$ for any nontrivial groups $A$ and $B$. Hence by Corollary 4.1.8 in [Bhagwat_2025], $L/K$ is primitive.
But the same proof doesn’t work for the case $r=n$ when $n$ has more than one prime in its prime factorization. In that case $L/K$ is Galois with $Gal(L/K)$ as $\mathbb{Z}/n\mathbb{Z}$. By chinese remainder theorem, $Gal(L/K)$ is isomorphic to $A\times B$ for some nontrivial groups $A$ and $B$ with $|A|>2$. Hence $L/K$ is not primitive similarly as in Remark 2.4.2.
@@ 2.4. Primitive Inverse Ascending Index Problem
For an extension $L/K$, we have the notion of ascending index of $L/K$, $t_{K}(L)$ which is $G:H^{G}]$ (where $H^{G}$ is normal closure of $H$ in $G$) and the quantity $u_{K}(L)$ which is $H^{G}:H]$ (See Section 7.2 and Section 9 in [Bhagwat_2025] for basic properties of ascending index of field extension).
By proof of Theorem 2.6 and by the proof of Inverse Ascending Index Problem for Number Fields Theorem 9.0.5 in [Bhagwat_2025], we have the following.
@@ 3. General Magnification
Just like the Cluster Magnification Theorem (Theorem 1 in [krithika2023root]) proved by the second and the third author inspired the first author and Bhagwat to define the concept of Strong Cluster Magnification and Weak Cluster Magnification in Section 4.1 and Section 4.3 in [Bhagwat_2025] respectively. In the same way, inspired from the Lemma 2 and the proof of the Main Theorem in [krithika2025inflated] by the second and the third author, we define in this section a more general notion of magnification. Then cluster magnification would turn out to be a special case of that.
@@ 3.1. Strong General Magnification
The following proposition justifies the terminology in Definition 3.1. The proof follows from the proof of Lemma 6.4 in [jaiswal2025rootcapacityintersectionindicium] but we still provide it for the sake of completeness.
@@
Since $\tilde{L}\cap\tilde{J}=K$. Hence $\tilde{L}\cap J^{\prime}=K$. Thus by Lemma 2.2.2 in [Bhagwat_2025] we have $\tilde{L}\cap L^{\prime}J^{\prime}=L^{\prime}$. Hence by Theorem 2.6 in [conrad2023galois], we have $Gal(\tilde{L}J^{\prime}/LJ^{\prime})\cong Gal(\tilde{L}/L)=H$ and $Gal(\tilde{L}J^{\prime}/L^{\prime}J^{\prime})\cong Gal(\tilde{L}/L^{\prime})=H_{0}$. Thus $LJ^{\prime}:L^{\prime}J^{\prime}]>1$ and $L_{1}\subset\tilde{L}J^{\prime}$. Similarly we have $L^{\prime}\cap\tilde{J}=K$ and hence $L^{\prime}J^{\prime}\cap\tilde{J}=J^{\prime}$ and $Gal(L^{\prime}\tilde{J}/L^{\prime}J)\cong Gal(\tilde{J}/J)=T$ and $Gal(L^{\prime}\tilde{J}/L^{\prime}J^{\prime})\cong Gal(\tilde{J}/J^{\prime})=T_{0}$ and $J_{1}\subset L^{\prime}\tilde{J}$.
Now $(\tilde{L}J^{\prime})(L^{\prime}\tilde{J})=\tilde{L}\tilde{J}=\tilde{M}$ and $Gal(\tilde{M}/L^{\prime}J^{\prime})=H_{0}\times T_{0}$. By Theorem 2.1 in [conrad2023galois] we have $\tilde{L}J^{\prime}\cap L^{\prime}\tilde{J}=L^{\prime}J^{\prime}$. Hence $L_{1}\cap J_{1}=L^{\prime}J^{\prime}$. Also $(LJ^{\prime})(L^{\prime}J)=LJ=M$. Finally since the Galois groups are isomorphic, we have that magnification tuple $(r_{L^{\prime}J^{\prime}}(L^{\prime}J),s_{L^{\prime}J^{\prime}}(L^{\prime}J),t_{L^{\prime}J^{\prime}}(L^{\prime}J),u_{L^{\prime}J^{\prime}}(L^{\prime}J))=(r_{J^{\prime}}(J),s_{J^{\prime}}(J),t_{J^{\prime}}(J),u_{J^{\prime}}(J))$.∎

### v2

# arXiv:2511.11841v2[math.NT] 28 Feb 2026

@@ Abstract.
In this article we establish certain variants of the Inverse Cluster Size problem. We introduce the notion of primitive extensions and establish the Primitive variant of the problem. Precisely, we prove the existence of primitive extensions over number fields of any given degree and cluster size less than the degree. We also introduce the notions of Strong and Weak General Magnification and the notion of general primitive extensions. We establish some interesting cases of the General primitive variant of the problem. We also recall the notion of totally real number fields and resolve the Totally real variant of the problem completely.
Key words and phrases:
inverse cluster size problem, primitive & general primitive extensions, totally real number fields
@@ 1. Introduction
In Galois theory, a basic result is that for a finite separable extension $L/K$ the number of automorphisms of $L$ fixing $K$ is equal to the degree $[L:K]$ if and only if $L/K$ is a Galois extension. In general, we have the notion of cluster size of $L/K$, $r_{K}(L)$ which is the number of automorphisms of $L$ fixing $K$ (See [1], [9], [12]). It is not hard to see that the cluster size will actually be a divisor of the degree $[L:K]$.
After proving Lagrange’s theorem in group theory, one naturally wonders about the converse and Cauchy’s theorem and Sylow’s theorem are such attempts at the converse. In the same spirit, a similar converse problem (which is called Inverse cluster size problem in [1]) can be asked : Given a perfect field $K$ and integers $n>2$ and $r\mid n$, does there exist an extension $L/K$ of degree $n$ with cluster size $r$ ?
The Inverse cluster size problem for $K=\mathbb{Q}$ was established in the pioneering work of Perlis [12], [11]. The Inverse cluster size problem for number fields was resolved completely by the first author and Bhagwat in [1] thus generalizing the result of Perlis for any number field. They used a similar approach using Shafarevich’s theorem and Hilbertian field theory along with certain Galois-theoretic lemmas which extended the result from $\mathbb{Q}$ to any number field $K$. This generalization also improved on the generalization proved previously by the second and third author in [9], which was proved by a process called Cluster magnification, reducing the problem to solving for the case $r=1$, but which excluded certain cases, namely $n=2r$, where $r$ is odd for $\mathbb{Q}$ and $n=2r$ for any number field $K\neq\mathbb{Q}$.
In this article we establish certain variants of the Inverse cluster size problem for number fields. In Section 2, the notion of an extension being primitive is introduced which means that the extension is not obtained from a nontrivial process of cluster magnification. We recall the notions of unique descending chains and unique ascending chains for extensions introduced in [1]. We then establish a criterion in Theorem 2.4 for an extension to be primitive in terms of unique chains which serves as a major ingredient in the proof of our Theorem 2.6 which is as follows.
@@
In Section 3, the notions of Strong and Weak General Magnification are introduced and studied inspired from the notions of Strong and Weak Cluster Magnification in [1] and motivated from results in [10]. In Section 4, the notion of an extension being general primitive is introduced which means that the extension is not obtained from a nontrivial process of general magnification. We then go on to establish some interesting cases of the General Primitive Inverse Cluster Size Problem which is encapsulated in the following result.
@@
In Section 5, we recall the notion of totally real number fields and resolve the Totally Real Variant of the problem completely which is as follows.
Theorem 1.3.
(Totally Real Inverse Cluster Size Problem)
Let $K$ be a totally real number field and $n>2$ and $r\mid n$. There exists an $L/K$ where $L$ is totally real such that $[L:K]=n$ and $r_{K}(L)=r$.
Furthermore we show existence of general primitive $L/K$ (hence also primitive) where $L$ is totally real number field for the following cases of $(K,n,r)$ in Theorem 1.2: Cases 1 (a), (b), (c) and Case $(K,2p,p)$ where $p>2$ is an odd prime.
Additionally we show existence of primitive $L/K$ where $L$ is totally real number field whenever we have $1<r<n$ for $n$ odd.
@@ 2. Primitive Inverse Cluster Size Problem
Let $K$ be a perfect field. We fix an algebraic closure $\bar{K}$ once and for all and work with finite extensions of $K$ contained in $\bar{K}$. Let $L/K$ be a finite extension and $\tilde{L}$ be its Galois closure inside $\bar{K}$. Let $G={{\rm Gal}}(\tilde{L}/K)$ and $H={{\rm Gal}}(\tilde{L}/L)$. We have the notion of cluster size of $L/K$, $r_{K}(L)$ which is the number of automorphisms of $L$ fixing $K$ which turns out to be the quantity $N_{G}(H):H]$ and thus a divisor of the degree $[L:K]=[G:H]$ (See Section 2.1 in [1]). We also have the notion of number of clusters of $L/K$, $s_{K}(L)$ which is $G:N_{G}(H)]$ which is also the number of distinct fields inside $\bar{K}$ isomorphic to $L$ over $K$ (See Section 3.2 in [1]).
@@ 2.1. Primitive Extensions
The following is reformulation of Definition 4.2.4 in [1] for field extensions (This notion for polynomials occurs in [9] as well).
@@
Consider $L/K$. Let $G={{\rm Gal}}(\tilde{L}/K)$ and $H={{\rm Gal}}(\tilde{L}/L)$. If $H$ is not contained in any proper normal subgroup of $G$ then $L/K$ is primitive. In particular, if $G$ is simple then $L/K$ is primitive.
@@ 2.2. Unique Chains for Extensions and Primitive Extensions
We have the notions of unique descending chains and unique ascending chains for extensions introduced in Section 7 in [1]. We recall those notions here. Let $L/K$ be a nontrivial finite extension.
@@
Let $\tilde{L}$ be Galois closure of $L/K$ and $G={{\rm Gal}}(\tilde{L}/K)$ and $H={{\rm Gal}}(\tilde{L}/L)$. By Theorem 7.1.1 in [1], $N_{1}=\tilde{L}^{N_{G}(H)}$. Now since $L=L^{\prime}F$ and $\tilde{L^{\prime}}\cap F=K$, we have that $L/L^{\prime}$ is Galois. Hence $L^{\prime}=\tilde{L}^{H^{\prime}}$ where $H\unlhd H^{\prime}\subset G$. Thus $H^{\prime}\subset N_{G}(H)$. Hence $N_{1}\subset L^{\prime}$. Thus $N_{i}\subset L^{\prime}$ and hence $F\subset L^{\prime}$. This forces $F=K$ which gives a contradiction to the magnification being nontrivial. ∎
@@
A shorter proof of above theorem by using results from Section 8.4 in [1]: We have $N_{i}=F_{j}\neq L,K$. Suppose $L/K$ is obtained by nontrivial strong cluster magnification from some subextension say $L^{\prime}/K$ through $F/K$. By Theorem 8.4.1 and Remark 8.4.2 in [1], $N_{i}\subset L^{\prime}$. By Theorem 8.4.3 and Remark 8.4.4 in [1], $F\subset F_{j}$. Thus $F\subset L^{\prime}$ which gives a contradiction.
@@
The assumption $N_{i}=F_{j}\neq L,K$ is important in the theorem. Suppose $L/K$ is Galois with Galois group $A\times B$ for $A$ and $B$ nontrivial groups and $|A|>2$. The unique descending chain for $L/K$ is $L=N_{0}\supsetneq K=N_{1}$ and unique ascending chain for $L/K$ is $K=F_{0}\subsetneq L=F_{1}$. Thus $N_{0}=F_{1}=L$ and $N_{1}=F_{0}=K$. But by Corollary 4.1.8 in [1] we have that $L/K$ is not primitive.
@@
The converse of Theorem 2.4 is false. Let $f$ over $K$ be irreducible of deg $n>2$ with Galois group ${\mathfrak{S}}_{n}$ with roots $\alpha_{i}\in\bar{K}$ for $1\leq i\leq n$. For $1\leq k\leq n-2$, let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$. Assume $k\neq 1,n/2$. By Theorem 7.3.6 in [1], the unique descending chain for $L_{k}/K$ is $L_{k}\supsetneq N_{k}$ and unique ascending chain for $L_{k}/K$ is $K$. Hence $L_{k}/K$ doesn’t satisfy hypothesis of Theorem 2.4 but we have that $L_{k}/K$ is primitive. This is because we don’t have any $K\subsetneq F\subset L_{k}$ such that $F/K$ is Galois.
@@
Since $c$ is odd, it does not divide the discriminant of $\mathbb{Q}(\zeta)/\mathbb{Q}$ where $\zeta$ is a primitive $n$-th root of unity in $\bar{\mathbb{Q}}$. Thus $\sqrt{c}\not\in\mathbb{Q}(\zeta)$. By Proposition 7.3.1 in [1], the unique descending chain is $L=N_{0}\supsetneq N_{1}\supsetneq\dots\supsetneq N_{k}=K$ and the unique ascending chain is $K=F_{0}\subsetneq F_{1}\subsetneq\dots\subsetneq F_{k}=L$ with $N_{i}=F_{k-i}=\mathbb{Q}(a^{2^{i}})$ for all $0\leq i\leq k$. Since $k\geq 2$, we are done by Theorem 2.4.
@@ 2.3. Primitive Inverse Cluster Size Problem
We proceed in the same way as the first author and Bhagwat proceeded in the proof of Inverse Cluster Size Problem for Number Fields Theorem 3.1.1 in [1]. Let $s=n/r$.
For $r=1$ we have that there exists $L/K$ of degree $n$ with ${{\rm Gal}}(\tilde{L}/K)$ as ${\mathfrak{S}}_{n}$. This $L/K$ satisfies $r_{K}(L)=1$. The subgroup of ${\mathfrak{S}}_{n}$ fixing $L$ is isomorphic to ${\mathfrak{S}}_{n-1}$. Since ${\mathfrak{S}}_{n-1}$ is not contained in any proper normal subgroup of ${\mathfrak{S}}_{n}$. Hence by Proposition 2.3, $L/K$ is primitive.
For $r>1$ we have that there exists $L/K$ with $G={{\rm Gal}}(\tilde{L}/K)$ solvable such that action of the group is transitive on $n$ points, and a point stabiliser fixes precisely $r$ points. We have $G=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}}$ with semidirect product group law
@@
Since $1<r<n$. Thus by Theorem 7.3.4 in [1], the unique descending chain is $L\supsetneq N\supsetneq K$ and the unique ascending chain is $K\subsetneq F\subsetneq L$ and they both coincide. Hence $N=F\neq L,K$. Thus by Theorem 2.4, $L/K$ is primitive. ∎
@@
The same proof works for the case $r=n=p^{k}>2$ where $p$ is a prime. In that case $L/K$ is Galois with ${{\rm Gal}}(L/K)$ as $\mathbb{Z}/p^{k}\mathbb{Z}$. By uniqueness in structure theorem for finite abelian groups, ${{\rm Gal}}(L/K)$ is not isomorphic to $A\times B$ for any nontrivial groups $A$ and $B$. Hence by Corollary 4.1.8 in [1], $L/K$ is primitive.
But the same proof doesn’t work for the case $r=n$ when $n$ has more than one prime in its prime factorization. In that case $L/K$ is Galois with ${{\rm Gal}}(L/K)$ as $\mathbb{Z}/n\mathbb{Z}$. By chinese remainder theorem, ${{\rm Gal}}(L/K)$ is isomorphic to $A\times B$ for some nontrivial groups $A$ and $B$ with $|A|>2$. Hence $L/K$ is not primitive similarly as in Remark 2.4.2.
@@ 2.4. Primitive Inverse Ascending Index Problem
For an extension $L/K$, we have the notion of ascending index of $L/K$, $t_{K}(L)$ which is $G:H^{G}]$ (where $H^{G}$ is normal closure of $H$ in $G$) and the quantity $u_{K}(L)$ which is $H^{G}:H]$ (See Section 7.2 and Section 9 in [1] for basic properties of ascending index of field extension).
By proof of Theorem 2.6 and by the proof of Inverse Ascending Index Problem for Number Fields Theorem 9.0.5 in [1], we have the following.
@@ 3. General Magnification
Just like the Cluster Magnification Theorem (Theorem 1 in [9]) proved by the second and the third author inspired the first author and Bhagwat to define the concept of Strong Cluster Magnification and Weak Cluster Magnification in Section 4.1 and Section 4.3 in [1] respectively. In the same way, inspired from the Lemma 2 and the proof of the Main Theorem in [10] by the second and the third author, we define in this section a more general notion of magnification. Then cluster magnification would turn out to be a special case of that.
@@ 3.1. Strong General Magnification
The following proposition justifies the terminology in Definition 3.1. The proof follows from the proof of Lemma 6.5 in [5] but we still provide it for the sake of completeness.
@@
Since $\tilde{L}\cap\tilde{J}=K$. Hence $\tilde{L}\cap J^{\prime}=K$. Thus by Lemma 2.2.2 in [1] we have $\tilde{L}\cap L^{\prime}J^{\prime}=L^{\prime}$. Hence by Theorem 2.6 in [4], we have ${{\rm Gal}}(\tilde{L}J^{\prime}/LJ^{\prime})\cong{{\rm Gal}}(\tilde{L}/L)=H$ and ${{\rm Gal}}(\tilde{L}J^{\prime}/L^{\prime}J^{\prime})\cong{{\rm Gal}}(\tilde{L}/L^{\prime})=H_{0}$. Thus $LJ^{\prime}:L^{\prime}J^{\prime}]>1$ and $L_{1}\subset\tilde{L}J^{\prime}$. Similarly we have $L^{\prime}\cap\tilde{J}=K$ and hence $L^{\prime}J^{\prime}\cap\tilde{J}=J^{\prime}$ and ${{\rm Gal}}(L^{\prime}\tilde{J}/L^{\prime}J)\cong{{\rm Gal}}(\tilde{J}/J)=T$ and ${{\rm Gal}}(L^{\prime}\tilde{J}/L^{\prime}J^{\prime})\cong{{\rm Gal}}(\tilde{J}/J^{\prime})=T_{0}$ and $J_{1}\subset L^{\prime}\tilde{J}$.
Now $(\tilde{L}J^{\prime})(L^{\prime}\tilde{J})=\tilde{L}\tilde{J}=\tilde{M}$ and ${{\rm Gal}}(\tilde{M}/L^{\prime}J^{\prime})=H_{0}\times T_{0}$. By Theorem 2.1 in [4] we have $\tilde{L}J^{\prime}\cap L^{\prime}\tilde{J}=L^{\prime}J^{\prime}$. Hence $L_{1}\cap J_{1}=L^{\prime}J^{\prime}$. Also $(LJ^{\prime})(L^{\prime}J)=LJ=M$. Finally since the Galois groups are isomorphic, we have that magnification tuple $(r_{L^{\prime}J^{\prime}}(L^{\prime}J),s_{L^{\prime}J^{\prime}}(L^{\prime}J),t_{L^{\prime}J^{\prime}}(L^{\prime}J),u_{L^{\prime}J^{\prime}}(L^{\prime}J))=(r_{J^{\prime}}(J),s_{J^{\prime}}(J),t_{J^{\prime}}(J),u_{J^{\prime}}(J))$.∎